MA483 Instructor: Russell Brown Spring 2001 Homework #3. Due Wednesday, 7 February 2001

• Strauss section 1.5, #1, 4.

• Strauss section 1.6, #5, 6 In #5, compute u_{xx} + 3u_{yy} − 2u_{x} + 24u_{y} + 5u when u has the special form u = e^{αx+βy}. Can you choose α and β to make the terms involving v_{x} and v_{y} go away? As a second step, make the change of variable x′ = x andy′ = γy so that in the new variable, the equation has the same coefficient in front of v_{x′x′} and v_{y′y′}.

• Strauss section 2.1 #3, 8 Announcements: Course handouts can be found at

http://www.ms.uky.edu/~rbrown/courses/ma483

The solutions to assigned homework are on reserve in the math library. Below are the e-mails of the students in this course.

• Ying Xu yxu0@engr.uky.edu
• Tapiwa Gurupira tzguru01@engr.uky.edu

• Pruittikorn Smithmaitre psmit0@engr.uky.edu

• Ravinder Siripuram ravi@engr.uky.edu

• James Booth jbooth@ms.uky.edu
• Delisa Nelson binky_4me@hotmail.com
• Mike Baker mbbakr@hotmail.com

January 31, 2001